# Enhancing Single Image to 3D Generation using Gaussian Splatting and Hybrid Diffusion Priors

### v1

@@ I INTRODUCTION
Two major unsolved challenges of 3D generation are visual fidelity (sharp and crisp outputs, well-defined textures) and geometric realism (geometrically consistent objects, that matches our real-life perception). Some recent efforts have fine-tuned existing 2D image diffusion models on 3D datasets, thereby enabling 3D geometric understanding [13, 7]. Foundation models are another promising direction [14, 15], enabling realistic 3D generation, thanks to large scale paired datasets [16] used for 2D-to-3D lifting. However, despite the ability of 3D-native methods to generate 3D-consistent assets—such as meshes, Gaussians, point clouds, or NeRF, sometimes within seconds, these methods continue to grapple with limitations in visual fidelity. Other works [17, 18] roughly combine 3D and 2D representations in spatial domain, blatently ignoring their distinctive outputs and geometric differences. This leads to lack of realism (e.g., assymetric structure, missing parts, etc.), and blurry reconstructions [19].
Unlike previous work, we propose combining 2D and 3D guidance in the frequency domain for two novel distillation losses, enabling both the textural and geometric guidance in our pipeline. Specifically, our contributions are as follows:

### v2

@@ I INTRODUCTION
Two major unsolved challenges of 3D generation are visual fidelity (sharp and crisp outputs, well-defined textures) and geometric realism (geometrically consistent objects, that match our real-life perception). Some recent efforts have fine-tuned existing 2D image diffusion models on 3D datasets, thereby enabling 3D geometric understanding [13, 7]. Foundation models are another promising direction [14, 15], enabling realistic 3D generation, thanks to large-scale paired datasets [16] used for 2D-to-3D lifting. However, despite the ability of 3D-native methods to generate 3D-consistent assets—such as meshes, Gaussians, point clouds, or NeRF, sometimes within seconds, these methods continue to grapple with limitations in visual fidelity. Other works [17] roughly combine 3D and 2D representations in the spatial domain, blatantly ignoring their distinctive outputs and geometric differences. This leads to a lack of realism (e.g., assymetric structure, missing parts, etc.), and blurry reconstructions [18].
Fourier123 [19], on the other hand, observed a discrepancy in the outputs of 2D and 3D prior-based models (Stable Diffusion [20] and Zero123 [13], respectively), both in image and frequency domains, and proposed a solution by proposing a 2D frequency-based SDS loss. However, they still fail to effectively combine the structural attributes from 2D priors and geometrical information from 3D prior, as visible in their findings. We draw motivation from this observation and propose a superior frequency-based combination of 2D and 3D priors. As highlighted in [21], geometric features can be attributed to low-frequency components whereas high-frequency components are majorly responsible for textures and edges. We, therefore propose two novel SDS loss functions utilizing 3D and 2D priors (details in IV-C). Instead of utilizing the entire frequency distribution (as in [19]), we filter the low and high-frequency components to selectively prioritize geometry and texture from 3D and 2D priors.
Unlike [19], we propose combining 2D and 3D guidance in the frequency domain for two novel distillation losses, enabling both the textural and geometric guidance in our pipeline. Specifically, our contributions are as follows:
@@ II-B Multi-view Diffusion for Mesh Reconstruction
Although our motivation aligns with Fourier123 [19], our methodology and contributions are fundamentally distinct: (1) Instead of relying on single-image Gaussian initialization, we utilize multi-view LGM [15] initialization, leveraging Zero123++ to generate novel views for a superior starting point; (2) Our method explicitly incorporates low-frequency components from 3D diffusion priors and high-frequency components from 2D diffusion priors—key elements absent from Fourier123’s approach; (3) We implement adaptive filtering, computing the cut-off frequency dynamically based on cumulative energy distribution in the amplitude spectrum, contrasting with Fourier123’s static amplitude-based SDS loss; (4) Sequential optimization in our framework surpasses the parallel optimization with hybrid SDS loss in Fourier123, yielding enhanced reconstruction, geometry, and texture fidelity; (5) Reference view guidance (Section IV-D) is integrated for additional supervision, a feature not addressed in Fourier123; and (6) Our work is specifically tailored for robotics, including dedicated experiments to adapt our approach for downstream tasks, representing a significant contribution to this field.